OpenStudy (anonymous):

y=x^2+8x-5.....It ask to type all intercepts....x=

OpenStudy (anonymous):

If you solve \[x^{2}+8x-5=0\]you will find the the x-intercepts.

OpenStudy (anonymous):

is the answer 8, -5

OpenStudy (anonymous):

oh, zero

OpenStudy (anonymous):

Ok, at this point i am tired and guessing. any help

OpenStudy (anonymous):

Sorry, I just feel really stupid right now.

OpenStudy (anonymous):

Ok, the quadratic formula will be of great use to you. It is \[x=\frac{ -b \pm \sqrt{b^{2}-4ac} }{ 2a }\]Where a is the coeffecient in front of x^2, b is the coefficient in front of x, and c is everything else.

OpenStudy (anonymous):

This is a good problem for the quadratic formula:\[x = \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\]where your equation is of the form: ax^2 + bx + c = 0 You have such an equation: x^2 + 8x - 5 = 0 So, a = 1, b = 8, and c = -5. Just substitute.

OpenStudy (anonymous):

You all seem so smart, I am older and never took algrbra

OpenStudy (anonymous):

Don't worry about age. Everyone can learn. The older student does have some advantages. Maturity, patience, perspective. You can do anything.

OpenStudy (anonymous):

so I would enter 1,8 and-5

OpenStudy (anonymous):

Yes. And don't get nervous or rush.

OpenStudy (anonymous):

Thanks for the confidence. I only have to more weeks and I have my degree, which was my goal because I put my children first with there education. Now, I feel I have waited to long. math is very hard for me.